1. Two objects are hung from strings. The top object (m1) has a mass of 30 kg and the bottom object (m2), which is hanging from m1, has a mass of 24 kg.

Calculate the tension in each string if you pull down on m1 with a force of 80 N [down].



1
Expert's answer
2021-03-17T16:53:23-0400


For "m_2"


"T_2-m_2g-F=0\to T_2=m_2g+F=24\cdot9.8+80=315.2\ (N)"

For "m_1"


"T_1-m_1g-T_2=0\to T_1=m_1g+T_2=30\cdot9.8+315.2=609.2\ (N)"